The homicide rate decreases at a rate of 5% per year in a city that had 900 homicides in the most recent year.
a. At this rate, in how many years will the number of homicides reach 450 in a year? Use the approximation formula.
B. What is the exact answer.

Answer by stanbon(75887)

The homicide rate decreases at a rate of 5% per year in a city that had 900 homicides in the most recent year.
a. At this rate, in how many years will the number of homicides reach 450 in a year? Use the approximation formula.
B. What is the exact answer.
450 = 900*0.95^t
0.95^t = 1/2
t = log(1/2)/log(0.95) = 13.51
---
Ans: In the 14th year
---------------
Cheers,
Stan H.
----------